### v1

# Non-adaptive Bellman-Ford:

Abstract
The Bellman-Ford algorithm for single-source shortest paths repeatedly updates tentative distances in an operation called relaxing an edge. In several important applications a non-adaptive (oblivious) implementation is preferred, which means fixing the entire sequence of relaxations upfront, independent of the edge-weights. In a dense graph on $n$ vertices, the algorithm in its standard form performs $(1+o(1))n^{3}$ relaxations. An improvement by Yen from 1970 reduces the number of relaxations by a factor of two. We show that no further constant-factor improvements are possible, and every non-adaptive deterministic algorithm based on relaxations must perform $(\frac{1}{2}-o(1))n^{3}$ steps. This improves an earlier lower bound of Eppstein of $(\frac{1}{6}-o(1))n^{3}$. Given that a non-adaptive randomized variant of Bellman-Ford with at most $(\frac{1}{3}+o(1))n^{3}$ relaxations (with high probability) is known, our result implies a strict separation between deterministic and randomized strategies, answering an open question of Eppstein.
1 Introduction
The Bellman-Ford algorithm [Bel58, For56, Moo59] is a classical method for computing single-source shortest paths in edge-weighted, directed graphs. Similarly to Dijkstra’s algorithm, the Bellman-Ford algorithm uses edge relaxations. Relaxing an edge $(u,v)$ means setting the tentative distance from the source $s$ to $v$ to the tentative distance from $s$ to $u$ plus the weight of the edge $(u,v)$, if this is smaller than the current tentative distance from $s$ to $v$. Initially setting all tentative distances to +$\infty$+ ∞, the algorithm proceeds via repeated edge-relaxations, updating tentative distances until they reach their final values, the true distances. Relaxing an edge $(u,v)$ changes the tentative distance from $s$ to $v$ to its final (correct) value, if and only if the tentative distance from $s$ to $u$ is already correct and $(u,v)$ is the last edge of a shortest path from $s$ to $v$. Algorithms of this type are also known as label-correcting.
Besides correctly handling negative edge-weights, an important feature of the Bellman-Ford algorithm is that it can be executed in a non-adaptive (also called oblivious) way. This means that the sequence of edge-relaxations is fixed in advance, independent of the edge-weights of the input graph and independent of the outcomes of previous operations. This reduces overhead while attaining the same asymptotic worst-case bound, and makes the algorithm more amenable for implementation in parallel or distributed settings, e.g., in network routing protocols [Hed88, MR17].
It is easy to see that a sequence of (non-adaptive) edge-relaxations correctly computes the distance from $s$ to $v$, if and only if the edges of some shortest path from $s$ to $v$ are relaxed in the order in which they appear on the path (possibly with other edge-relaxations in-between). Moreover, if for every possible path from $s$ to $v$, the edges of the path are relaxed in the order they appear going from $s$ to $v$, then the correct distance from $s$ to $v$ must have been computed. Conversely, if the edges of some path do not appear as a subsequence of the relaxation sequence, then, for some configuration of the edge-weights (e.g., with weights on the path set to 00 and all other weights set to $1$) the correct distance will not have been reached.
The problem of finding the most efficient edge-relaxation sequence thus maps to the combinatorial problem of finding the shortest sequence of edges in which every possible path from $s$ to another vertex appears as a subsequence. The standard version of the Bellman-Ford algorithm relaxes edges in a round-robin fashion. In its non-adaptive implementation, this requires relaxing each edge $n-1$ times, where $n$ is the number of vertices. In a complete directed graph this leads to $(1+o(1))n^{3}$ relaxations and a proportional overall running time.
A classical improvement by Yen [Yen70] reduces the number of relaxations by a factor of two, while preserving the non-adaptive nature of the algorithm. Yen’s algorithm decomposes the edge-set of the input graph into forward edges and backward edges, both sets forming acyclic subgraphs. It then alternates between relaxing all forward- and all backward edges in topological order.
The question of whether a better relaxation order exists than the one given by Yen’s method has been open since 1970. Recently, Eppstein [Epp23] showed that every non-adaptive deterministic algorithm must do $(\frac{1}{6}+o(1))n^{3}$ relaxations in dense graphs, leaving open the possibility of a further factor-three improvement. Bannister and Eppstein [BE12] obtained a randomized improvement to Yen’s algorithm, with a high-probability guarantee on the algorithm’s success. In the non-adaptive setting, this improves Yen’s bound by a factor of $\frac{2}{3}$.
When adaptivity is allowed, i.e., the algorithm can decide which edge to relax based on the outcomes of past relaxations (and possibly other calculations involving edge-weights), better bounds are known (e.g., see [Epp23] and references therein), this setting is, however, not our concern in this paper. Lower bounds for adaptive algorithms have so far been elusive.
Our result.
We show that every non-adaptive deterministic algorithm based on edge-relaxations must perform $(\frac{1}{2}-o(1))n^{3}$ relaxation steps on a complete directed graph with $n$ vertices. This shows that Yen’s algorithm cannot be improved with the choice of a better relaxation sequence, by any constant factor. Moreover, since faster randomized approaches are known, randomization strictly helps for non-adaptive relaxation sequences. This answers an open question of Eppstein [Epp23].
2 Deterministic lower bound
Let $G$ be the complete directed graph with vertex set $V$ and edge set $E$, with $|V|=n$ and $|E|=n(n-1)$, and let $s\in V$ be a special vertex (the source). Let $w:E\rightarrow\mathbb{R}$ denote the edge-weights. Note that $G$ is simple, containing no loops or multiple edges.
The Bellman-Ford algorithm and its variants compute the distances from $s$ to every vertex in $V$ by initially setting $d(s,s)=0$ and $d(s,v)=+\infty$ for all $v\in V\setminus\{s\}$, and repeatedly performing operations $\operatorname{\mathsf{relax}}(e)$ for edges $e\in E$. The operation $\operatorname{\mathsf{relax}}(e)$ for $e=(u,v)$ sets $d(s,v)\leftarrow\min\{d(s,v),d(s,u)+w(e)\}$. The goal is to execute such steps until $d(s,v)$ is the correct distance from $s$ to $v$ for all $v\in V$. We assume that the graph contains no negative weight cycle (such a cycle can be detected by checking if any $\operatorname{\mathsf{relax}}$ operation still reduces tentative distances after the algorithm has finished).
As we consider non-adaptive algorithms, the behaviour of the algorithm is fully specified by the sequence of relax operations $S=(e_{1},\dots,e_{L})$ it performs, where $e_{i}\in E$ for all $1\leq i\leq L$. We refer to $L=|S|$ as the length of the sequence.
All paths considered are simple. We say that a sequence $S$ of edges contains a path $(v_{1},\dots,v_{k})$ if $S$ contains $(v_{1},v_{2}),\dots,(v_{k-1},v_{k})$ as a subsequence (in the given order, but not necessarily contiguously). A $k$-path is a path $(s,v_{1},\dots,v_{k})$ for some set of $k$ distinct vertices $v_{1},\dots,v_{k}\in V\setminus\{s\}$. We say that a sequence of edges $S$ is $k$-valid, if it contains every possible $k$-path. By our earlier discussion, a sequence $S$ is a valid relaxation sequence in a non-adaptive single-source shortest path algorithm, if and only if it is $(n-1)$-valid.
Let $T_{k}$ denote the length of the shortest $k$-valid sequence for graph $G$.
Theorem 2.1.
$T_{k}\geq{(k-2)}{{n-1}\choose{2}}+(n-1)^{2}$, for all $12\leq k\leq n-$ 1.
Our claimed result follows by setting $k=n-1$ and observing $(n-3){{n-1}\choose{2}}+(n-1)^{2}\geq\frac{n^{3}}{2}-O(n^{2})$. In the remainder of the section we prove Theorem 2.1.
We proceed by induction on $k$. For $k=2$, the claim holds: to contain every $2$-path, a sequence must contain every edge except those pointing to $s$, and thus have length at least $(n-1)^{2}$.
Assume now $k>2$ and let $S=(e_{1},\dots,e_{T_{k}})$ be a shortest $k$-valid sequence. Observe that if $S$ contains every $k$-path, then it also contains every $(k-1)$-path, i.e., it is $(k-1)$-valid.
For each entry $e_{i}$ of $S$, call $e_{i}$ critical, if and only if some $h$-path $P$ with $h\leq k-1$ is contained in $(e_{1},\dots,e_{i})$ but not in $(e_{1},\dots,e_{i-1})$. In words, $e_{i}$ finishes $P$ and is the first edge in the sequence to do so. Otherwise, call $e_{i}$ non-critical.
Let $N$ be the number of non-critical entries in $S$. Observe that $T_{k}=|S|\geq T_{k-1}+N$. This is because, as the non-critical edges are not needed for finishing any prefix of a $(k-1)$-path, the subsequence of $S$ obtained by removing all non-critical edges is still $(k-1)$-valid, and thus of length at least $T_{k-1}$.
Let $q_{i}$ be the endpoint of $e_{i}$, i.e., $e_{i}=(x,q_{i})$, for some $x,q_{i}\in V$. Let $Q_{i}$ denote the set of endpoints of critical edges in the sequence $(e_{i},\dots,e_{T_{k}})$. More precisely, set $Q_{T_{k}+1}=\emptyset$, and for $i=T_{k},\dots,1$, set $Q_{i}=Q_{i+1}\cup\{q_{i}\}$, if $e_{i}$ is critical, and set $Q_{i}=Q_{i+1}$ if $e_{i}$ is non-critical. Let us also denote, $\bar{Q}_{i}=(V\setminus\{s\})\setminus Q_{i}$.
Call a critical entry $e_{i}$ supercritical whenever $|Q_{i}|>|Q_{i+1}|$. In words, such an edge $e_{i}$ adds the new endpoint $q_{i}$ to the set of critical endpoints as we go right to left. As $S$ contains paths ending in every vertex (other than the source), notice that $Q_{1}=V\setminus\{s\}$. This means that there are $n-1$ supercritical entries in $S$. Denote these (left-to-right) by $e_{i_{1}},\dots,e_{i_{n-1}}$, with their (distinct) endpoints $q_{i_{1}},\dots,q_{i_{n-1}}$.
[…]
Repeating the argument for each supercritical edge, we conclude that all edges $(q_{i_{b}},q_{i_{a}})$ for $11\leq a<b\leq n-$ 1 must appear as non-critical entries. This implies a lower bound on the number of non-critical entries $N\geq{{n-1}\choose{2}}$, from which $T_{k}=|S|\geq T_{k-1}+{{n-1}\choose{2}}$ follows. Using the induction hypothesis to lower bound $T_{k-1}$ yields the claimed result.

### v2

@@
 The Bellman-Ford algorithm for single-source shortest paths repeatedly updates tentative distances in an operation called relaxing an edge.
-In several important applications a non-adaptive (oblivious) implementation is preferred, which means fixing the entire sequence of relaxations upfront, independent of the edge-weights.
-In a dense graph on $n$ vertices, the algorithm in its standard form performs $(1+o(1))n^{3}$ relaxations.
+In several important applications a non-adaptive (oblivious) implementation is preferred, which means fixing the entire sequence of relaxations upfront, independently of the edge-weights.
+Such an implementation performs, in a dense graph on $n$ vertices, $(1+o(1))n^{3}$ relaxations.
 An improvement by Yen from 1970 reduces the number of relaxations by a factor of two.
@@
 Given that a non-adaptive randomized variant of Bellman-Ford with at most $(\frac{1}{3}+o(1))n^{3}$ relaxations (with high probability) is known, our result implies a strict separation between deterministic and randomized strategies, answering an open question of Eppstein.
+On the complexity side, we show that deciding whether a given relaxation sequence is guaranteed to yield correct distances is NP-hard, even with the complete graph as input.
 1 Introduction
@@
 Similarly to Dijkstra’s algorithm, the Bellman-Ford algorithm uses edge relaxations.
-Relaxing an edge $(u,v)$ means setting the tentative distance from the source $s$ to $v$ to the tentative distance from $s$ to $u$ plus the weight of the edge $(u,v)$, if this is smaller than the current tentative distance from $s$ to $v$.
-Initially setting all tentative distances to +$\infty$+ ∞, the algorithm proceeds via repeated edge-relaxations, updating tentative distances until they reach their final values, the true distances.
+Relaxing an edge $(u,v)$ means setting the tentative distance from the source $s$ to $v$ to the tentative distances from $s$ to $u$ plus the weight of the edge $(u,v)$, if this is smaller than the current tentative distance from $s$ to $v$.
+Initially setting the tentative distance from $s$ to $s$ itself to 00 and to all other vertices to +$\infty$+ ∞, the algorithm proceeds via repeated edge-relaxations, updating tentative distances until they reach their final values, the true distances.
 Relaxing an edge $(u,v)$ changes the tentative distance from $s$ to $v$ to its final (correct) value, if and only if the tentative distance from $s$ to $u$ is already correct and $(u,v)$ is the last edge of a shortest path from $s$ to $v$.
 Algorithms of this type are also known as label-correcting.
-Besides correctly handling negative edge-weights, an important feature of the Bellman-Ford algorithm is that it can be executed in a non-adaptive (also called oblivious) way.
-This means that the sequence of edge-relaxations is fixed in advance, independent of the edge-weights of the input graph and independent of the outcomes of previous operations.
-This reduces overhead while attaining the same asymptotic worst-case bound, and makes the algorithm more amenable for implementation in parallel or distributed settings, e.g., in network routing protocols [Hed88, MR17].
-It is easy to see that a sequence of (non-adaptive) edge-relaxations correctly computes the distance from $s$ to $v$, if and only if the edges of some shortest path from $s$ to $v$ are relaxed in the order in which they appear on the path (possibly with other edge-relaxations in-between).
-Moreover, if for every possible path from $s$ to $v$, the edges of the path are relaxed in the order they appear going from $s$ to $v$, then the correct distance from $s$ to $v$ must have been computed.
-Conversely, if the edges of some path do not appear as a subsequence of the relaxation sequence, then, for some configuration of the edge-weights (e.g., with weights on the path set to 00 and all other weights set to $1$) the correct distance will not have been reached.
-The problem of finding the most efficient edge-relaxation sequence thus maps to the combinatorial problem of finding the shortest sequence of edges in which every possible path from $s$ to another vertex appears as a subsequence.
-The standard version of the Bellman-Ford algorithm relaxes edges in a round-robin fashion.
-In its non-adaptive implementation, this requires relaxing each edge $n-1$ times, where $n$ is the number of vertices.
-In a complete directed graph this leads to $(1+o(1))n^{3}$ relaxations and a proportional overall running time.
+Besides the correct handling of negative edge-weights, an important feature of the Bellman-Ford algorithm is that it can be executed in a non-adaptive (also called oblivious) way.
+This means that the sequence of edge-relaxations is fixed in advance, independently of the edge-weights of the input graph and of the outcomes of previous operations.
+This reduces overhead while maintaining the same asymptotic worst-case running time.
+A non-adaptive sequence of edge-relaxations also makes the algorithm amenable for implementation in parallel or distributed settings, e.g., in network routing protocols [Hed88, MR17], where centralized control may not be available.
+It is easy to see that a sequence of (non-adaptive) edge-relaxations correctly computes the distance from $s$ to $v$ if and only if the edges of some shortest path from $s$ to $v$ are relaxed in the order in which they appear on the path (possibly with other edge-relaxations in between).
+If for every simple path from $s$ to $v$, the edges of the path are relaxed in the order they appear going from $s$ to $v$, then the correct distance from $s$ to $v$ must have been computed.
+Conversely, if the edges of some simple $s$-to-$v$ path do not appear as a subsequence of the relaxation sequence, then, for some configuration of the edge-weights (e.g., with weights on the path set to 00 and all other weights set to $1$) the correct distance will not have been reached.
+The problem of finding the most efficient edge-relaxation sequence thus maps to the combinatorial problem of finding the shortest sequence of edges in which every simple path from $s$ to another vertex appears as a subsequence.
+The standard non-adaptive version of the Bellman-Ford algorithm relaxes all $m$ edges in a round-robin fashion $n-1$ times, where $n$ is the number of vertices.
+The resulting sequence of edges clearly contains every simple path, and in a complete directed graph the number of edge-relaxations is $(1+o(1))n^{3}$, with a proportional overall running time.
 A classical improvement by Yen [Yen70] reduces the number of relaxations by a factor of two, while preserving the non-adaptive nature of the algorithm.
-Yen’s algorithm decomposes the edge-set of the input graph into forward edges and backward edges, both sets forming acyclic subgraphs.
+Yen’s algorithm decomposes the edge set of the input graph into forward edges and backward edges, both sets forming acyclic subgraphs.
 It then alternates between relaxing all forward- and all backward edges in topological order.
 The question of whether a better relaxation order exists than the one given by Yen’s method has been open since 1970.
-Recently, Eppstein [Epp23] showed that every non-adaptive deterministic algorithm must do $(\frac{1}{6}+o(1))n^{3}$ relaxations in dense graphs, leaving open the possibility of a further factor-three improvement.
+Recently, Eppstein [Epp23] showed that every non-adaptive deterministic algorithm must perform $(\frac{1}{6}-o(1))n^{3}$ relaxations in dense graphs, leaving open the possibility of a further factor-three improvement.
 Bannister and Eppstein [BE12] obtained a randomized improvement to Yen’s algorithm, with a high-probability guarantee on the algorithm’s success.
 In the non-adaptive setting, this improves Yen’s bound by a factor of $\frac{2}{3}$.
-When adaptivity is allowed, i.e., the algorithm can decide which edge to relax based on the outcomes of past relaxations (and possibly other calculations involving edge-weights), better bounds are known (e.g., see [Epp23] and references therein), this setting is, however, not our concern in this paper.
-Lower bounds for adaptive algorithms have so far been elusive.
-Our result.
-We show that every non-adaptive deterministic algorithm based on edge-relaxations must perform $(\frac{1}{2}-o(1))n^{3}$ relaxation steps on a complete directed graph with $n$ vertices.
+Our results.
+We show (§ 2) that every non-adaptive deterministic algorithm based on edge-relaxations must perform $(\frac{1}{2}-o(1))n^{3}$ relaxation steps on a complete directed graph with $n$ vertices.
 This shows that Yen’s algorithm cannot be improved with the choice of a better relaxation sequence, by any constant factor.
@@
 This answers an open question of Eppstein [Epp23].
+On the complexity side, two natural questions arise.
+Given a directed graph $G$ and a source vertex $s$, we may ask for a shortest non-adaptive sequence of edge-relaxations that compute the correct distances from $s$, for all possible configurations of edge-weights.
+In light of the previous discussion, this amounts to finding a shortest sequence of edges that contains as subsequence every simple path from $s$ in $G$.
+A second question is to decide, whether a given sequence $S$ of edges is a valid relaxation sequence, i.e., whether it contains every simple path from $s$ in $G$.
+While we suspect both problems to be (NP-)hard, we can only show this for the second, and leave the first one open. The hardness of the second problem holds even for the complete graph $G$ (§ 3).
+We note that both problems admit polynomial-time algorithms in the easier case of BF-orderable graphs, i.e., graphs in which every edge needs to be relaxed at most once; this results from 1980s work of Mehlhorn and Schmidt [MS86], and Haddad and Schäffer [HS88].
+Further related work.
+Due to the fundamental nature of the single-source shortest paths problem in the presence of negative weights, several attempts were made to establish its precise complexity or to improve the long-standing $O(mn)$ runtime of the Bellman-Ford algorithm.
+We refer to Eppstein [Epp23] for a discussion of results in different models and for variants of the problem.
+Note that complexity bounds are sensitive to the exact modeling assumptions, in particular, they differ between the non-adaptive and adaptive settings.
+When adaptivity is allowed, i.e., the algorithm can decide which edge to relax based on the outcomes of past relaxations (and possibly other calculations involving edge-weights), a recent breakthrough by Fineman [Fin24] has reduced the complexity to $\tilde{O}(mn^{8/9})$, with a later improvement [HJQ25] to $\tilde{O}(mn^{4/5})$.
+These results may explain why lower bounds for adaptive algorithms have been elusive.
+Nevertheless, an asymptotic lower bound of $\Omega(n^{3})$ has recently been shown [AHS+24], for adaptive relaxation-based algorithms with a certain restricted set of operations (while classical Bellman-Ford-variants fall within these restrictions, the mentioned recent improvements do not).
 2 Deterministic lower bound
-Let $G$ be the complete directed graph with vertex set $V$ and edge set $E$, with $|V|=n$ and $|E|=n(n-1)$, and let $s\in V$ be a special vertex (the source).
+Let $G$ be the complete directed graph with vertex set $V$ and edge set $E$, with $|V|=n$ and $|E|=n(n-1)$, and let $s\in V$ be the source vertex.
 Let $w:E\rightarrow\mathbb{R}$ denote the edge-weights.
@@
 The goal is to execute such steps until $d(s,v)$ is the correct distance from $s$ to $v$ for all $v\in V$.
-We assume that the graph contains no negative weight cycle (such a cycle can be detected by checking if any $\operatorname{\mathsf{relax}}$ operation still reduces tentative distances after the algorithm has finished).
+We assume that the graph contains no negative weight cycle (such a cycle can be detected by checking if any relax operation still reduces tentative distances after the algorithm has finished).
 As we consider non-adaptive algorithms, the behaviour of the algorithm is fully specified by the sequence of relax operations $S=(e_{1},\dots,e_{L})$ it performs, where $e_{i}\in E$ for all $1\leq i\leq L$.
 We refer to $L=|S|$ as the length of the sequence.
+Our main result is the following.
+Theorem 2.1.
+Every valid relaxation sequence for $G$ must have length at least $(\frac{1}{2}-o(1))n^{3}$.
+In the remainder of the section we prove the theorem.
 All paths considered are simple.
-We say that a sequence $S$ of edges contains a path $(v_{1},\dots,v_{k})$ if $S$ contains $(v_{1},v_{2}),\dots,(v_{k-1},v_{k})$ as a subsequence (in the given order, but not necessarily contiguously).
-A $k$-path is a path $(s,v_{1},\dots,v_{k})$ for some set of $k$ distinct vertices $v_{1},\dots,v_{k}\in V\setminus\{s\}$.
+We say that a sequence $S$ of edges contains a path $(v_{1},\dots,v_{k})$ if $S$ contains $(v_{1},v_{2}),(v_{2},v_{3}),\dots,(v_{k-1},v_{k})$ as a subsequence (in the given order, but not necessarily contiguously).
+A $k$-path is a path $(s,v_{1},\dots,v_{k})$ for some set of $k$ vertices $\{v_{1},\dots,v_{k}\}\subseteq V\setminus\{s\}$.
 We say that a sequence of edges $S$ is $k$-valid, if it contains every possible $k$-path.
-By our earlier discussion, a sequence $S$ is a valid relaxation sequence in a non-adaptive single-source shortest path algorithm, if and only if it is $(n-1)$-valid.
-Let $T_{k}$ denote the length of the shortest $k$-valid sequence for graph $G$.
+By our earlier discussion, a sequence $S$ is a valid relaxation sequence for $G$ in a non-adaptive single-source shortest path algorithm, if and only if it is $(n-1)$-valid.
+The task of finding a short $k$-valid sequence appears related to the task of finding a short permutation supersequence, a well-studied question in combinatorics.
+In that context, $c_{n}$ denotes the smallest possible length of a sequence over $\{1,\dots,n\}$ that contains each permutation of $\{1,\dots,n\}$ as a subsequence (not necessarily contiguously).
+It is known that $c_{n}\leq\lceil n^{2}-\frac{7}{3}n+\frac{19}{3}\rceil$ [Rad12], see also the survey of Engen and Vatter [EV20].
+Most relevantly for us, a lower bound of the form $c_{n}\geq(1-o(1))n^{2}$ was shown in 1976 by Kleitman and Kwiatkowski.
+Lemma 2.2 ([KK76]).
+For all $n\geq 1$ and all $\varepsilon>0$, we have $c_{n}>n^{2}-C_{\varepsilon}\cdot n^{7/4+\varepsilon}$, with $C_{\varepsilon}$ depending only on $\varepsilon$.
+We will use Lemma 2.2 to prove Theorem 2.1.111In an earlier version of the paper we gave a more direct proof of Theorem 2.1, but the proof was faulty.
+It is well-known that whenever $N$ is even, the edge-set of a complete (undirected) graph on $N$ vertices can be decomposed into $N-1$ disjoint perfect matchings, i.e., sets of $\frac{N}{2}$ non-incident edges.
+(Such a decomposition is called a $1$-factorization, often studied in the context of round-robin tournaments, e.g., see [MR85]).
+If $N$ is odd, we can simply ignore one of the vertices and decompose the remainder of the graph into $N-2$ disjoint matchings, each consisting of $\frac{N-1}{2}$ edges.
+Denote the vertex set of the complete directed graph $G$ as $V=\{s,v_{1},\dots,v_{n-1}\}$.
+Consider a complete (undirected) graph on $\{v_{1},\dots,v_{n-1}\}$.
+By the previous paragraph, regardless of the parity of $n$, we can find in this graph $n-3$ disjoint matchings, each of size $\frac{n-2}{2}$.
+Denote such a collection of matchings as $E_{1},\dots,E_{n-3}$ (each $E_{i}$ is a set of undirected edges).
+For each matching $E_{i}$ we form two sets of directed edges in $G$; one where each edge of $E_{i}$ is oriented from smaller to larger index vertex (call this $F_{i}$), and one where each edge of $E_{i}$ is oriented from larger to smaller index vertex (call this ${H_{i}}$).
+We thus obtain a collection =$\mathcal{F}=\{F_{1},H_{1},\dots,F_{n-3},{H_{n-3}}\}$ of pairwise disjoint sets, each consisting of pairwise non-incident directed edges in $G$.
+Let $S$ be a shortest $(n-1)$-valid, or simply valid, sequence of edges for graph $G$. For any $F\in\mathcal{F}$ let $S_{F}$ denote the maximal subsequence of $S$ consisting of elements in $F$.
+We claim that $S_{F}$ must contain every permutation of $F$ as a subsequence, and thus the lower bound of Lemma 2.2 applies to $S_{F}$.
+Indeed, suppose some permutation of $F$ is not contained in $S_{F}$, and denote this permutation (of directed edges of $G$) as $(e_{1},\dots,e_{\frac{n-2}{2}})$, where $e_{i}$ is the directed edge $(x_{i},y_{i})$, for $x_{i},y_{i}\in V\setminus\{s\}$.
+In this case, $S$ could not contain the sequence of edges $((s,x_{1}),\mathbf{e_{1}},(y_{1},x_{2}),\mathbf{e_{2}},(y_{2},x_{3}),\dots,\mathbf{e_{\frac{n-2}{2}}})$ that form a simple path from the source to $y_{\frac{n-2}{2}}$, and thus $S$ could not be valid.
+Since $|F|=\frac{n-2}{2}$ for all $F\in\mathcal{F}$, we have $|S_{F}|\geq(\frac{1}{4}-o(1))n^{2}$, by Lemma 2.2.
+As edge sets in $\mathcal{F}$ are pairwise disjoint, the corresponding subsequences of $S$ do not overlap.
+Thus, we have
+$|S|~{}\geq~{}|\mathcal{F}|\cdot\left(\frac{1}{4}-o(1)\right)n^{2}~{}\geq~{}\left(2n-6\right)\left(\frac{1}{4}-o(1)\right)n^{2}~{}\geq~{}\left(\frac{1}{2}-o(1)\right)n^{3}.$
+This completes the proof. $\hfill\qed$
+3 Hardness of deciding validity
+Again, let $G$ be a directed graph, with a special vertex $s$.
+Let $S=(e_{1},\dots,e_{|S|})$ be a sequence of edges of $G$.
+The relaxation validity question asks whether $S$ contains as subsequence every simple path of $G$ starting in $s$, in other words, whether $S$ is a valid relaxation sequence.
 Theorem 3.1.
-$T_{k}\geq{(k-2)}{{n-1}\choose{2}}+(n-1)^{2}$, for all $12\leq k\leq n-$ 1.
-Our claimed result follows by setting $k=n-1$ and observing $(n-3){{n-1}\choose{2}}+(n-1)^{2}\geq\frac{n^{3}}{2}-O(n^{2})$. In the remainder of the section we prove Theorem 2.1.
-We proceed by induction on $k$. For $k=2$, the claim holds: to contain every $2$-path, a sequence must contain every edge except those pointing to $s$, and thus have length at least $(n-1)^{2}$.
-Assume now $k>2$ and let $S=(e_{1},\dots,e_{T_{k}})$ be a shortest $k$-valid sequence. Observe that if $S$ contains every $k$-path, then it also contains every $(k-1)$-path, i.e., it is $(k-1)$-valid.
-For each entry $e_{i}$ of $S$, call $e_{i}$ critical, if and only if some $h$-path $P$ with $h\leq k-1$ is contained in $(e_{1},\dots,e_{i})$ but not in $(e_{1},\dots,e_{i-1})$.
-In words, $e_{i}$ finishes $P$ and is the first edge in the sequence to do so.
-Otherwise, call $e_{i}$ non-critical.
-Let $N$ be the number of non-critical entries in $S$.
-Observe that $T_{k}=|S|\geq T_{k-1}+N$.
-This is because, as the non-critical edges are not needed for finishing any prefix of a $(k-1)$-path, the subsequence of $S$ obtained by removing all non-critical edges is still $(k-1)$-valid, and thus of length at least $T_{k-1}$.
-Let $q_{i}$ be the endpoint of $e_{i}$, i.e., $e_{i}=(x,q_{i})$, for some $x,q_{i}\in V$.
-Let $Q_{i}$ denote the set of endpoints of critical edges in the sequence $(e_{i},\dots,e_{T_{k}})$.
-More precisely, set $Q_{T_{k}+1}=\emptyset$, and for $i=T_{k},\dots,1$, set $Q_{i}=Q_{i+1}\cup\{q_{i}\}$, if $e_{i}$ is critical, and set $Q_{i}=Q_{i+1}$ if $e_{i}$ is non-critical.
-Let us also denote, $\bar{Q}_{i}=(V\setminus\{s\})\setminus Q_{i}$.
-Call a critical entry $e_{i}$ supercritical whenever $|Q_{i}|>|Q_{i+1}|$.
-In words, such an edge $e_{i}$ adds the new endpoint $q_{i}$ to the set of critical endpoints as we go right to left.
-As $S$ contains paths ending in every vertex (other than the source), notice that $Q_{1}=V\setminus\{s\}$.
-This means that there are $n-1$ supercritical entries in $S$.
-Denote these (left-to-right) by $e_{i_{1}},\dots,e_{i_{n-1}}$, with their (distinct) endpoints $q_{i_{1}},\dots,q_{i_{n-1}}$.
-[…]
-Repeating the argument for each supercritical edge, we conclude that all edges $(q_{i_{b}},q_{i_{a}})$ for $11\leq a<b\leq n-$ 1 must appear as non-critical entries.
-This implies a lower bound on the number of non-critical entries $N\geq{{n-1}\choose{2}}$, from which $T_{k}=|S|\geq T_{k-1}+{{n-1}\choose{2}}$ follows.
-Using the induction hypothesis to lower bound $T_{k-1}$ yields the claimed result.
+Relaxation validity is co-NP-complete.
+In the remainder of the section we prove Theorem 3.1.
+Membership in co-NP is trivial; if $S$ is not valid, then some simple path $P$ witnesses this.
+Verifying that $S$ does not contain $P$ can be done by greedily matching edges of $P$ in their order from $s$, to the earliest occurrence in $S$, to the right of the entry matched to the predecessor edge.
+To show co-NP-hardness, we reduce from the all-permutations supersequence (APS) problem, shown to be co-NP-hard by Uznański [Uzn15].
+APS asks whether a given sequence $T=(t_{1},\dots,t_{|T|})$ contains every permutation of $\{1,\dots,n\}$ as a subsequence.
+(Alternatively, one may ask whether there is a permutation not contained in $T$, which is consequently NP-hard).
+The reduction from APS to relaxation validity is straightforward.
+Consider a complete directed graph $G$ with a set of $12n+$ 1 vertices denoted $V=\{s,x_{1},y_{1},\dots,x_{n},y_{n}\}$.
+For $i\in\{1,\dots,n\}$, let $e_{i}$ denote the special edge $(x_{i},y_{i})$.
+Let $Q$ be a valid relaxation sequence of $G$, i.e., a $2n$-valid sequence of edges. Let $R$ be the sequence of edges obtained from $Q$ by deleting every occurrence of $e_{i}$ for all $i$.
+We construct a relaxation sequence $S$ of $G$ as follows.
+For $i=1,\dots,|T|$, we form $S_{i}$ by appending to the end of sequence $R$ the edge $e_{j}$, for $j=t_{i}$.
+Then, we concatenate $S_{1},\dots,S_{|T|}$, and finally $R$, to form the sequence $S$.
+From the preceding discussion, we can enforce $|Q|,|R|\in O(n^{3})$ and the construction of $S$ takes time polynomial in $|T|$.
+(Notice that we do not require $Q$ to be optimal, a simple round-robin sequence of length $O(n^{3})$ is sufficient).
+We claim that $T$ contains every permutation of $\{1,\dots,n\}$ if and only if $S$ contains every simple path starting from $s$ in $G$, from which Theorem 3.1 follows.
+First, suppose $T$ does not contain some permutation $\pi_{1},\dots,\pi_{n}$.
+Then, $S$ does not contain the sequence of edges $(e_{\pi_{1}},\dots,e_{\pi_{n}})$ as these edges only appear in $S$ in correspondence to entries of $T$.
+Consequently, $S$ cannot correctly relax the path $(s,x_{\pi_{1}},y_{\pi_{1}},\dots,x_{\pi_{n}},y_{\pi_{n}})$.
+Suppose now that $T$ contains every permutation of $\{1,\dots,n\}$.
+Then, every simple $k$-path $P$ is correctly relaxed by $S$.
+Indeed, since $T$ contains every permutation, $S$ contains the (possibly incomplete) permutation of edges $e_{\pi_{1}},e_{\pi_{2}},\dots$ in $P$, in their order of appearance in $P$.
+The portions of $P$ between any two edges $e_{\pi_{i}}$, $e_{\pi_{i+1}}$ or before the first, respectively, after the last such edge, are necessarily contained in some copy of $R$ inserted between the corresponding occurrences of $e_{\pi_{i}}$, $e_{\pi_{i+1}}$ in $S$, resp., at the beginning or at the end of $S$.
+This concludes the proof.
+4 Conclusions
+We showed that the number of edge-relaxations in Yen’s variant of the Bellman-Ford algorithm cannot be improved in a deterministic non-adaptive way by any constant factor.
+Our lower bound is shown for the complete directed graph as the input.
+Besides serving as a worst-case example, the complete graph is also natural as it captures full non-adaptivity: If an algorithm is oblivious to both the weights and the structure of the graph, then it must attempt to relax every possible edge, as if the input were a complete graph.
+In the randomized case, a gap between the $(\frac{1}{12}-o(1))n^{3}$ lower bound due to Eppstein and the $(\frac{1}{3}+o(1))n^{3}$ upper bound due to Bannister and Eppstein remains [Epp23, BE12].
+For incomplete dense graphs, i.e., with the number of edges $m\geq n^{1+{\varepsilon}}$, Eppstein shows that in the worst case, the $O(mn)$ bound cannot be improved by non-adaptive relaxation sequences (in this case, the algorithm knows the graph structure but not the weights).
+For sparse families, a gap between the $O(mn)$ upper bound and an $\Omega(mn/\log{n})$ lower bound remains [Epp23].
+The complexity of finding a short relaxation sequence for a given graph remains open.
+The problem appears similar to the shortest supersequence problem, known to be NP-hard in several variants (e.g., see [Mai78, RU81]).
+The difficulty of extending these results to relaxation sequences is twofold.
+The paths in our setting are simple, i.e., without repeated vertices or edges, whereas in the inputs to the shortest supersequence problem symbols can appear multiple times.
+To our knowledge, the complexity of finding the shortest supersequence of a set of (partial) permutations is open.
+A second difficulty is that the set of input paths in our setting is only implicitly given and the number of paths is, in general, exponential.